OpenStudy (anonymous):

I Will give medals ! (: Explain, in complete sentences, how you would divide the following expression: 10√x7 divided by 7√x4 Write the simplified answer in radical form.

OpenStudy (pokemon23):

Hero can solve that with his eyes open

OpenStudy (mathmate):

The trick is put everything under radical signs under one single radical sign, and simplify. 10√x7 divided by 7√x4 =(10/7)* √(x^7/x^4) =(10/7)√x^3 =(10x√x)/7

hero (hero):

I thought you wanted that explained using complete sentences